OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of three over four division symbol seven over twelve? four-thirds times twelve-sevenths three-fourths times twelve-sevenths four-thirds times seven-twelfths seven-twelfths times three-fourths

OpenStudy (psymon):

Before I just work through the problem, do you actually want me to go step by step? I know the answer was just given to you last time, so I rather not just explain if you would prefer to just have the answer.

OpenStudy (anonymous):

Step By Step is better. :/

OpenStudy (psymon):

Alright. I'm only asking, often people just want answers and don't wanna listen. So when you divide fractions, you actually need to take the 2nd fraction and flip it upside down. They call this the reciprocal, but it's really just flipping the fraction. So your problem is this: \[\frac{ 3 }{ 4 }\div \frac{ 7 }{ 12 }\]That being said, to solve this we want to flip that 7/12 upside down. Once we flip that fraction, we multiply straight across. So this problem will become this: \[\frac{ 3 }{ 4 }*\frac{ 12 }{ 7 }\]Now just like before, we would multiply straight across. The 3 and 12 on top multiply and the 4 multiplies with 7 on bottom. There is a slightly easier way to do this, but I just want to make sure we know the basic concept first. Make sense?

OpenStudy (anonymous):

Yeah

OpenStudy (psymon):

Okay, cool :3 So we can just multiply acrosstoget \[\frac{ 3*12 }{ 4*7 } = \frac{ 36 }{ 28 }\] Now we just need to reduce this into lowest terms. Are you comfortable doing that part?

OpenStudy (anonymous):

Um...

OpenStudy (anonymous):

No lol

OpenStudy (psymon):

Lol, alright, no problem. SO what we want to do is find a number that will divide evenly into the top and bottom. The numbers we choose must be the same. If we divide the top by 2 we must divide the bottom by 2. Sometimes you may not see what the largest number that divides into both top and bottom is, but that's fine. If you're not sure what number to choose, this is something you might try. Let's break apart 36 and 28. We will find all the factors (numbers that divide evenly into it) of 36 and 28 and see what the largest number they have in common is. So if I start by breaking apart 36 I would get: 36| 2, 18, 3, 12, 4, 9, 6 Now if I do this to 28: 28| 2, 14, 4, 7 From this you can see that the biggest number they have in common is 4. So this will be the number I choose to divide the top and bottom by. I will do it like this: \[\frac{ 36\div4 }{ 28\div4 } = \frac{ 9 }{ 7 }\] So this would be the answer to the division problem. Now I know I typed out a lot, so make sure you understand what I did.

OpenStudy (anonymous):

But 9 /7 Isnt a Option

OpenStudy (psymon):

Well your question just wants the first step. The first step is what I did earlier by flipping the 2nd fraction and then beginning to multiply. I just worked us out through the answer. So in terms of just the first step. it would be (3/4)*(12/7). I'm just saying that we still would need to know how to actually solve it.

OpenStudy (anonymous):

OH

OpenStudy (psymon):

Mhm. I justfinished theproblem instead of stopping. So the main idea is doyou see how I tried to reduce it?